### v1

# Single-Source Regular Path Querying in Terms of Linear Algebra

Abstract.
A given edge-labelled graph two-way regular path queries (2-RPQs) allow one to use regular languages over labelled edges and inverted edges to constraint paths of interest. 2-RPQs are (partially) adopted in different real-world graph analysis systems and are a part of the GQL ISO standard. But the performance of 2-RPQs on real-world graphs is still a bottleneck for wider adoption. A new single-source 2-RPQ algorithm based on linear algebra is proposed. Utilization of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries. Our implementation demonstrates better performance on average on Wikidata and the respective query log in comparison with MillenniumDB, FalkorDB, and the algorithm of Diego Arroyuelo et al.
Graph Querying, Regular Path Querying, RPQ, Graph Analysis Algorithms, Sparse Linear Algebra, GraphBLAS
copyright: acmlicensedjournalyear: 2024conference: ; ; booktitle: ccs: Information systems Graph-based database modelsccs: Information systems Query languages for non-relational enginesccs: Theory of computation Regular languagesccs: Mathematics of computing Paths and connectivity problems
1. Introduction
Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are formulated in terms of a formal language. The language restricts the set of accepted paths: the sentence formed by the labels of a path should be in the language. Regular languages are the most popular class of constraints used as navigational queries in graph databases.
Queries that use regular languages to specify constraints are called regular path queries or RPQs. Being proposed in 1989 by Alberto O. Mendelzon and Peter T. Wood (Mendelzon and Wood, 1989), RPQs studying intensively for now. Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extend the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b). Regular path constraints and related extensions adopted (partially, in some cases) in a wide range of graph analysis systems and respective query languages, such as Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016). Moreover, RPQ is a part of ISO standard for a graph query language GQL [^1] ISO/IEC 39075:2024, GQL, https://www.iso.org/standard/76120.html, and core of SPARQL 1.1 RDF query language (201, 2013).
Nevertheless, having a long history of theoretical and applied research, adoption in real-world solutions, RPQs (and related extensions) are still in the focus of research. One of the important direction is an implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real-world cases, and performance-targeted solutions still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024). Thus, the development of new performant algorithms for RPQs evaluation is an actual challenge that also pointed out by Angela Bonifati et al in (Bonifati et al., 2018). Set of different solutions was proposed to improve RPQ evaluation performance: from special indexes creation (Liu et al., 2021; Arroyuelo et al., 2023b) to utilization of different parallel and distributed models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
Sparse linear algebra is a promising way to high-performance graph analysis. This idea is actively developed by GraphBLAS (Kepner et al., 2016) community [^2] GraphBLAS community web page: https://graphblas.org/. Huge amount of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^3] Almost full list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers. and respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021). Even more recently, it was shown that utilization of sparse linear algebra is a way to a high-performance algorithm for more expressive class of queries, Context-Free Path Queries, or CFPQ (Terekhov et al., 2021). But, to our knowledge, there are only a few results on linear algebra based algorithms for RPQ. On the one hand, modern graph database FalkorBD444 Sources of FalkorBD on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and utilises linear algebra to analyse graphs. This database supports a subset of the Cypher graph query language, including parts of regular constraints. But there is no detailed analysis of respective algorithm. At the same time, linear algebra based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a). The proposed solution directly translates regular expressions to expressions over matrices and vectors. Being implemented using sparse matrices and parallel computations, this solution still has performance problems that are evident from the evaluation (Arroyuelo et al., 2023a).
There are BFS-based strategies for RPQ evaluation (e.g. (Koschmieder and Leser, 2012)) [^5] For overview of the primary techniques to evaluate regular path queries, such as relational algebra based and automata based approaches, and ways to optimise it, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018).. At the same time, BFS and its variations can be expressed in terms of linear algebra (Davis, 2019a). It opens a way to develop a linear algebra based algorithm for RPQ evaluation having BFS-like traversal as the core, and we investigate this way.
To summarise, in this work we make the following contributions.
  1. (1)
    A new BFS-inspired algorithm (LA 2-RPQ) for single-source (or single-destination as a symmetric case) 2-RPQ is proposed. The algorithm is based on linear algebra: it is expressed in terms of operations over sparse boolean matrices. This fact allows one to utilise high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation. The correctness of the proposed algorithm is proven.
  2. (2)
    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), and with the state-of-the-art graph database MillenniumDB. As a result, we show that our solution is comparable with competitors on real-world dataset Wikidata and the respective query log proposed in MillenniumDB path query challenge (Farias et al., 2023). While for some queries the proposed algorithm is slower than the competitors, it is faster on average: ×$\times 3.9$ for MillenniumDB, ×$\times 19.6$ for FalkorDB, ×$\times 4.2$ for the algorithm of Diego Arroyuelo et al. Moreover, it fits within the 1 minute time limit for all queries, while the other competitors require more time for some queries.
2. Preliminaries
In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution.
2.1. Basics of graph theory
First we define the graph that we use as a data model.
Definition 2.0 (Edge-labelled graph).
A quadruple
$\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is called an edge-labelled graph (or a graph) if:
  • $V$ is a finite set of vertices;
  • $E\subseteq V\times V$ is a finite set of edges;
  • $L$ is a finite set of labels;
  • $\lambda_{\mathcal{G}}:E\mapsto 2^{L}$ represents edge labels.
Any finite set can be enumerated by natural numbers from $1$ to $n$. For the rest of the paper, we will assume that the vertices of the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ are enumerated, and without loss of generality $V=\{1,2,...,|V|\}$.
Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an edge-labelled graph. Introduce symbols and sets:
  • $a^{-}\notin L$ for $a\in L$, $(a^{-})^{-}=a$ and $a=b\Leftrightarrow a^{-}=b^{-}$
  • $e^{-}=(v,u)$ where $e=(u,v)$.
  • $E^{-}=\{e^{-}\mid e\in E\}$.
  • $L^{-}=\{a^{-}\mid a\in L\}$.
  • $\lambda_{\mathcal{G}}^{-}:E^{-}\mapsto 2^{L^{-}},\;\lambda_{\mathcal{G}}^{-}(e^{-})=\{a^{-}\mid a\in\lambda_{\mathcal{G}}(e)\}$.
We also need these sets to generalize the definition of the directed graph to be able to traverse it in both directions:
  • $E^{\leftrightarrow}=E\cup E^{-}$.
  • $L^{\leftrightarrow}=L\cup L^{-}$.
  • $\lambda_{\mathcal{G}}^{\leftrightarrow}:E^{\leftrightarrow}\mapsto 2^{L^{\leftrightarrow}},\;\;\lambda_{\mathcal{G}}^{\leftrightarrow}(e)=\lambda_{\mathcal{G}}(e)\cup\lambda^{-}_{\mathcal{G}}(e)\;$.
  • $\mathcal{G}^{\leftrightarrow}=\langle V,E^{\leftrightarrow},L^{\leftrightarrow},\lambda^{\leftrightarrow}\rangle$.
Definition 2.0 (path).
A path $\pi=(e_{1},e_{2},...,e_{n})$ of length $n$ in the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ from $u_{1}$ to $v_{n}$ is a finite sequence of edges $e_{i}=(u_{i},v_{i})\in E$ s.t. $\forall 1\leq i\leq n-1\;v_{i}=u_{i+1}$.
We say there are zero-length paths represented by an empty sequence from $v$ to $v$ for all $v\in V$.
Definition 2.0 (2-way path).
A 2-way path (2-path) $\pi=(e_{1},e_{2},...,e_{n})$ in the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a path in $\mathcal{G}^{\leftrightarrow}=\langle V,E^{\leftrightarrow},L^{\leftrightarrow},\lambda_{\mathcal{G}}^{\leftrightarrow}\rangle$.
The $\omega$ maps paths to words as defined below.
  • $\omega_{\mathcal{G}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{G}}(e_{i})\}$ for path $\pi=(e_{1},e_{2},...,e_{n})$ in $\mathcal{G}$.
  • $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda^{\leftrightarrow}_{\mathcal{G}}(e_{i})\}$ for 2-path $\pi=(e_{1},e_{2},...,e_{n})$ in $\mathcal{G}$.
where $\cdot$ denotes concatenation.
Regular languages (RLs) represent the set of all languages that are accepted by finite automata. However, we are going to look for 2-way paths for which we need an NFA modification to be introduced.
Definition 2.0 (2-way non-deterministic finite automaton).
A 2-way non-deterministic finite automaton (2-NFA) is a tuple $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$, where:
  • $Q$ is a finite set of states;
  • $\Sigma$ is a finite alphabet;
  • $\Delta\subseteq Q\times Q$ is transition relation;
  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ maps transitions to labels;
  • $Q_{S}\subseteq Q$ is a set of starting states;
  • $Q_{F}\subseteq Q$ is a set of final states.
The set of languages accepted by 2-NFAs coincides with the set of all RLs over $\Sigma^{\leftrightarrow}$. Let $[\![\mathcal{N}]\!]=\mathcal{R}$ where $\mathcal{R}$ is the RL accepted by the 2-NFA $\mathcal{N}$.
In practice special kinds of strings called regular expressions provide a useful facility to define regular languages. The algorithms supplying the transformations between regular expressions, RLs and 2-NFAs are out of the scope of this paper.
Notice that 2-NFA can be seen as a graph $\mathcal{G}_{\mathcal{N}}=\langle Q,\Delta,\Sigma^{\leftrightarrow},\lambda_{\mathcal{N}}\rangle$ equipped with the set of starting states $Q_{S}\subseteq Q$ and the set of final states $Q_{F}\subseteq Q$. Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
Vice versa a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can bee seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$. Such 2-NFA accepts the language:
$[\![\mathcal{N}_{G}]\!]=\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})$
Thus, we can treat evaluation of 2-RPQs with fixed set of starting vertices as an intersection of 2-NFAs.
Definition 2.0 (2-NFA intersection).
For two arbitrary 2-NFAs $\mathcal{N}_{1}=\langle Q_{1},\Sigma_{1},\Delta_{1},\lambda_{\mathcal{N}_{1}},Q_{S1},Q_{F1}\rangle$ and $\mathcal{N}_{2}=\langle Q_{2},\Sigma_{2},\Delta_{2},\lambda_{\mathcal{N}_{2}},Q_{S2},Q_{F2}\rangle$ introduce a new automaton $\mathcal{N}=\mathcal{N}_{1}\times\mathcal{N}_{2}$ called the intersection of 2-NFAs $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ where $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ s.t.:
  • $Q=Q_{1}\times Q_{2}$;
  • $\Sigma=\Sigma_{1}\cap\Sigma_{2}$;
  • $\Delta=\{((q_{1},q_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\mid(q_{1},q_{1}^{\prime})\in\Delta_{1}$, $(q_{2},q_{2}^{\prime})\in\Delta_{2}\}$;
  • $\lambda_{\mathcal{N}}(((q_{1},q_{2}),(q_{1}^{\prime},q_{2}^{\prime})))=\lambda_{\mathcal{N}_{1}}((q_{1},q_{1}^{\prime}))\cap\lambda_{\mathcal{N}_{2}}((q_{2},q_{2}^{\prime}))$;
  • $Q_{S}=Q_{S1}\times Q_{S2}$;
  • $Q_{F}=Q_{F1}\times Q_{F2}$.
The resulting 2-NFA accepts the intersection of the languages defined by the initial automata $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$:
$[\![\mathcal{N}_{1}\times\mathcal{N}_{2}]\!]=[\![\mathcal{N}_{1}]\!]\cap[\![\mathcal{N}_{2}]\!].$
We have all the required preliminaries to formally state the problem solved by the 2-RPQ algorithm.
Definition 2.0 (Two-way regular path query).
Recall two-way regular path query (2-RPQ) a triple $\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$ and $v_{s}\in V$ is a starting node.
Frequent practical cases are queries with fixed starting or final vertex. So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ we are interested in an efficient way of evaluating these maps:
  • Single-source reachability:
    $[\![\mathcal{Q}]\!]_{SSR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }u\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
  • Single-destination reachability:
    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{s}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
3. Linear Algebra, Graphs and Relations
In order to operate with graphs in terms of linear algebra we need to see how algebraic object are connected with set relations and how the relations are connected to graph theory.
For two enumerated finite sets $A=\{1,2,...n\}$, $B=\{1,2,...,m\}$ for some $n,m\in\mathbf{N}$ a binary matrix $T$ of size $|A|\times|B|$ can be used to represent a relation $\mathcal{T}\subseteq A\times B$: $T_{ij}=1\Leftrightarrow(i,j)\in\mathcal{T}$. Recall this matrix $T$ a matrix representing the relation $\mathcal{T}$.
We need to define the following linear algebra operations over the Boolean matrices. Let $A$, $B$, $C$ be the matrices over Boolean algebra =$\langle\mathcal{B}=\{0,1\},\vee,\wedge,\neg,0,1\rangle$. Introduce the following operations and constants.
  • Sum: $A_{n\times m}\operatorname{\operatorname{\oplus}}B_{n\times m}=(a_{ij}\vee b_{ij})_{n\times m}$.
  • Product: $A_{n\times m}\operatorname{\operatorname{\otimes}}B_{m\times k}=\displaystyle\left(\;\bigvee_{k=1}^{n}a_{ik}\wedge b_{kj}\;\right)_{n\times k}$.
  • Transposition: $A_{n\times m}^{T}=(a_{ji})_{m\times n}$.
  • Zero-matrix: $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}=(0)_{n\times k}$.
  • Mask: $A_{n\times m}\langle C_{n\times m}\rangle=(a_{ij}\wedge c_{ij})_{n\times m}$
  • Complement: $\neg A=(\neg a_{ij})$
Let $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$ be binary relations over sets $S_{1}$, $S_{2}$, $S_{3}$. Let $A$, $B$ and $C$ be binary matrices representing them. Then the following correspondences hold.
  • $C=A\operatorname{\operatorname{\oplus}}B$ $\leftrightarrow\mathcal{C}=\mathcal{A}\cup\mathcal{B}$
           where $\mathcal{A},\mathcal{B},\mathcal{C}\subseteq S_{1}\times S_{2}$.
  • $C=A\operatorname{\operatorname{\otimes}}B$ $\leftrightarrow\mathcal{C}=\{(i,k)\mid(i,j)\in\mathcal{A}$, $(j,k)\in\mathcal{B}\}$
           if $\mathcal{A}\subseteq S_{1}\times S_{2}$, $\mathcal{B}\subseteq S_{2}\times S_{3}$ and
           $\mathcal{C}\subseteq S_{1}\times S_{3}$.
  • $C=A^{T}$ $\leftrightarrow\mathcal{C}=\{(i,j)\mid(j,i)\in\mathcal{A}\}=\mathcal{A}^{-}$
           where $\mathcal{A}\subseteq S_{1}\times S_{2}$, $C\subseteq S_{2}\times S_{1}$.
  • $C=A\langle B\rangle$ $\leftrightarrow\mathcal{C}=\mathcal{A}\cap\mathcal{B}$,
           where $\mathcal{A},\mathcal{B},\mathcal{C}\subseteq S_{1}\times S_{2}$.
  • $C=\neg A$ $\leftrightarrow\mathcal{C}=\{(i,j)\mid(i,j)\in S_{1}\times S_{2}\setminus\mathcal{A}\}$
           where $\mathcal{A},\mathcal{C}\subseteq S_{1}\times S_{2}$.
Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an arbitrary graph and $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ an arbitrary 2-NFA. Introduce the sets:
  • $E^{a}=\{e\mid e\in E,a\in\lambda^{\leftrightarrow}_{\mathcal{G}}(e)\}$ for all $a\in L^{\leftrightarrow}$.
  • $\Delta^{a}=\{\delta\mid\delta\in\Delta,a\in\lambda_{\mathcal{N}}(\delta)\}$ for all $a\in\Sigma^{\leftrightarrow}$.
These sets consist of the edges of the graph $\mathcal{G}$ and the transitions of 2-NFA $\mathcal{N}$ marked with a label $a$. They can be seen as binary relations over the sets $V$ and $Q$ correspondingly.
Definition 3.0 (Adjacency matrix of the label).
Let $G^{a}$ be the matrix representing the binary relation $E^{a}$. $G^{a}$ is called an adjacency matrix of the label $a$.
Definition 3.0 (Boolean decomposition of the adjacency matrix).
A Boolean decomposition of the adjacency matrix $G$ is a set of Boolean matrices $\{G^{a}\mid a\in L\}$.
Boolean decomposition of the adjacency matrices of some real-world data represented by graphs is a set of sparse matrices. This fact strictly leads to the idea of exploiting it for an efficient representation and using sparse matrix operation algorithm implementations.
Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{-}$ and $G^{a^{-}}=(G^{a})^{T}$.
4. Single-Source RPQ in Terms of Linear Algebra
In this section we describe a single-source linear-algebra based 2-RPQ algorithm and prove its correctness.
input : $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}}Q_{S},Q_{F}\rangle$, $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $v_{s}\in V$
output : Matrix $P_{|Q|\times|V|}$
1
2$\{N^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{N}$ adjacency matrix;
3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
4 $M_{|Q|\times|V|},P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
5 for $q\in Q_{S}$ do
       $M_{q{v_{s}}}\leftarrow 1$;
        // Set the starting relation
6
7 end for
8while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
       $M\leftarrow\bigoplus_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}((N^{a})^{T}\operatorname{\operatorname{\otimes}}M\operatorname{\operatorname{\otimes}}G^{a})\langle\neg P\rangle$;
        // Update $\mathcal{M}$
       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$ ;
        // Accumulate resulting matrix
9
10 end while
11
return $P$
Algorithm 1 Single Source 2-RPQ using linear algebra
Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA represents query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$. Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be the input graph and $v_{s}\in V$ is a starting node. Then the algorithm 1 builds the following automata intersection:
$[\![\mathcal{N}\times\mathcal{N}_{G}]\!]=[\![\mathcal{N}]\!]\cap[\![\mathcal{N}_{\mathcal{G}}]\!]=\mathcal{R}\cap\left(\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})\right).$
Core of the algorithm is a line 8 that performs one step of traversing two automata simultaneously. Result of the algorithm is a matrix $P_{|Q|\times|V|}$ such that
$\begin{cases}P_{q_{F}v}=1\\ q_{F}\in Q_{F}\end{cases}\Leftrightarrow\begin{cases}\exists\pi_{\mathcal{G}}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\mathcal{R}\neq\operatorname{\varnothing}.\end{cases}$
An example of such a step is presented in figure 1. For graph and automaton with labels $\{a,b\}$, we visualise adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation. Solid-colored edges should be traversed simultaneously at the step. Dashed edges used to represent relation $M$: the edge connects two vertex that are simultaneously reachable. Initially, such an edge connects the initial state of the automaton and the starting vertex in the graph. After the step (one iteration of the main cycle), there are two such edges. One of them connects the final state of the automaton with vertex 5 of the graph. That means vertex 5 is reachable from the start vertex 3 by the path that farms a word acceptable by the automaton.
Hence, the set of reachable nodes can be easily obtained from the matrix $P$. Let $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ be an evaluated 2-RPQ. Then by the main property of the algorithm 1 resulting matrix $P$ is such that
$v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P_{q_{F}v}=1\text{ for some }q_{F}\in Q_{F}.$
Procedure 2 computes the explicit representation of the query result. Namely, $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$. Formally, the following theorem holds.
Theorem 4.1 (LA 2-RPQ algorithm correctness).
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
$(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
input : Matrix $P_{|Q|\times|V|}$, set of final states $Q_{F}$
output : Matrix $P^{F}_{1\times|V|}$
1
[…]
3 for $q\in Q_{F}$ do
       $F_{1q_{F}}\leftarrow 1$;
        // Setting up a matrix reducing $P$
4
5 end for
return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
Algorithm 2 Exporting reachable vertices subprogram
Proof of this theorem is straightforward induction by the paths length. Details are provided in appendix A.
Particularly, from the theorem, we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$. As far as $q_{F}$ is a final state of the NFA, $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\subseteq[\![\mathcal{N}]\!]$. Thus, $v$ reachable from $v_{s}$ by path satisfies the given regular constraint.
Figure 1. The algorithm step for graph and automaton for regular expression $a^{*}b$
The proposed single-source 2-RPQ algorithm can be used to solve the single-destination problem. Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in single destination query. Denote:
  • =$\mathcal{R}^{-}=\{w^{-}=a_{n}^{-}a_{n-1}^{-}...a_{2}^{-}a_{1}^{-}\mid w=a_{1}a_{2}...a_{n})\in\mathcal{R}\}$
  • $\mathcal{N}^{-}=\langle Q,\Sigma,\Delta^{-},\lambda_{\mathcal{N}}^{-},Q_{F},Q_{S}\rangle$.
  • $\mathcal{Q}^{-}=\langle\mathcal{G},\mathcal{R}^{-},v_{s}\rangle$.
The automaton $\mathcal{N}^{-}$ allows traversing paths in the opposite direction. It holds $\mathcal{R}^{-}=[\![\mathcal{N}^{-}]\!]$. Evaluating single-source reachability 2-RPQ on the reversed 2-RPQ $\mathcal{Q}^{-}$ gives using the 2-NFA $\mathcal{N}^{-}$:
$[\![\mathcal{Q}]\!]_{SDR}=[\![\mathcal{Q}^{-}]\!]_{SSR}.$
Boolean matrix decomposition of $\mathcal{N}^{-}$ can be easily obtained by transposing the matrices in Boolean matrix decomposition of $\mathcal{N}$.
For performance reasons, it’s necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer non-zero entries. Associativity of the Boolean matrix multiplication can be used to perform less calculation. This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1:
$((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
In general, the first option is preferable due to the fact that 2-NFA for the query can be converted into the corresponding minimal deterministic finite automaton having no more than $|Q|$ non-zero cells in each $\mathcal{N}^{a}$ adjacency matrix due to the automaton determinism. This strictly leads to less expensive computation of the product with large $|V|\times|V|$ matrix.
5. Evaluation
The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) inside the LAGraph (Szárnyas et al., 2021) infrastructure [^6] The source code of the algorithm is available in our GitHub LAGraph fork: https://github.com/SparseLinearAlgebra/LAGraph/tree/rpq. Both sparse matrices and their transpositions are loaded in-memory due to the fact we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
Whereas we’re focusing on an efficient way of evaluating RPQs in-memory, there aren’t many candidates to compare with. Popular database management systems primarily focus on general availability and ensure availability of the concurrent access at the same time the suggested algorithm implementation is suited only for solving the reachability problem. We’ve selected the following systems as demonstrating the best performance and being the most related to our use case.
MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system supporting a synthetic way to carry out calculations without using the disk storage by caching the query data. It demonstrates state-of-the-art performance on evaluating regular path queries.
FalkorDB (ex. RedisGraph (Cailliau et al., 2019)) is an in-memory property-graph database also employing SuiteSparse:GraphBLAS for query evaluation. However, it uses OpenCypher modification and supports only a subset of regular path queries. To deal with this, we’ve carried out the measurements of cypher-compatible and non-cypher-compatible queries separately (e.g., it is impossible to evaluate repeated-path queries in the form of $(ab)^{*}$ in FalkorDB).
RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the in-memory linear-algebra based RPQ evaluation algorithm. The original work introduces a few variations of the adjacency matrix representation: $k^{2}$-trees offering less memory consumption and CSR/CSC formats providing better performance. We’ve chosen the last one to compare with since we’re aiming to compare the performance.
All experiments are conducted on a work station with Ryzen 9 7900X 4.7 GHz 12-core, 128 Gb of DDR5 RAM and running Ubuntu 22.04.
5.1. Dataset Description
For algorithm evaluation, we choose the Wikidata dataset from the snapshot provided in terms of MillenniumDB path query challenge (Farias et al., 2023). The resource contains both the graph and the set of 660 different 2-RPQs in SPARQL format taken from the Wikidata query log.
For systems that do not support RDF format, the dataset was converted to an edge-labelled graph. The processed graph has 610 402 396610402396610\,402\,396610 402 396 edges, 91 609 2549160925491\,609\,25491 609 254 vertices and 1 39513951\,3951 395 distinct labels.
SPARQL path queries has been converted to the corresponding minimized DFAs. For evaluating on FalkorDB the supported subset of regular path queries have been converted into Cypher queries of form MATCH ... COUNT (DISTINCT ...). The original dataset contains queries from all to all vertices and queries involving non-existing entities. They’ve been filtered out, resulting in 578 queries remaining.
5.2. Evaluation Results
Figure 2. Wikidata dataset per-query evaluation time
LA 2-RPQ MilenniumDB RPQ-Matrix FalkorDB
Mean C (ms) 83 327 348 1 625
Median C (ms) 2.5 4.4 4.2 18.8
Mean speedup C 1 3.9 4.2 19.6
Median speedup C 1 1.8 1.7 7.5
Mean NC (ms) 374 1 345 7 108
Median NC (ms) 37.4 69.8 269.6
Mean speedup NC 1 3.6 19.0
Median speedup NC 1 1.9 7.2
Memory (Gb) 18.1 6.9 50.0
BPT 31.9 12.1 85.8
Timeouts 0 1 15 53
Table 1. Query execution results. Results are provided for cypher-compatible (C) and non-cypher-compatible (NC) queries separately. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
We measure evaluation time per query with a timeout of 1 minute. All of the data required for query evaluation is pre-loaded into the memory. Queries are executed one by one.
We run MillenniumDB [^7] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 in a SPARQL mode with the maximum amount of CPU cores available. The query dataset is executed twice. The first run isn’t measured since it makes the database load the data into the memory. The query execution time is calculated as the internal optimizer plus execution time excluding the parser.
We execute FalkorDB v4.2.0888FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 querying through the python wrapper. The database is started without CPU core limitations and without query cache. We build an index over the vertices to be able to select the starting or destination vertex using it. The query execution time is based on the internal timer.
The RPQ-matrix algorithm is executed in CSR/CSC mode without core and memory consumption limitations. The consumed time is measured with the internal timer. The solution has been built out of the project repository [^9] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606.
Results of the measurements, both evaluation time and relative speedup, are provided in figure 2. Only queries expressible in Cypher are presented. The purple lines represent mean values, and the green lines represent medians. Note that relative speedup is represented as pointwise speedup for each query. So, mean and median calculated for per-query speedup. Numeric values of the mean and the median, the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are available within the table 1. Queries that have failed on any of the systems with timeouts are excluded from the statistics. The number of time outed queries shown in table 1.
We can see that while in median, speeding up for our solution is between 1.7 and 7.5, in mean we improve performance more than 19 times (relative to the most closest solutions RPQ-Matrix and FalkorDB). That means the proposed solution significantly outperforms other linear algebra based solutions on some queries (up to ×$\times 10^{5}$) while demonstrating comparable (and better in average) performance in the most cases. On the other hand, RPQ-Matrix demonstrates better performance on big number of queries. It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based solutions (proposed and RPQ-Matrix) to take the best from both of them.
Note that MillenniumDB is pretty stable: there is neither outstanding wins nor losses in performance. At the same time, the proposed solution is slightly better in average even though MillenniumDB caches query evaluation data, but our solution does not.
Memory consumption is presented on table 1. We don’t specify MillenniumDB memory consumption due to the fact it can process query in-memory only by caching data and doesn’t provide ways to preload whole dataset from the disk storage. RPQ-Matrix demonstrates the lowest memory usage. Memory consumption of the proposed solution is comparable with competitors.
We collected the top 10 queries with the worst and best speedup relative to all competitors. Collected data presented in tables 345. We can see that the weakness of our solution are relatively simple queries with a relatively small number of reachable vertices, like Q P131* ?x1 with less than 100 reachable vertices, or ?x1 (P138 / P106) Q with less than a million reachable vertices, where Q is a fixed vertex, and ?x1 is a variable. The forte of the proposed solution is a complex query (in terms of regexp complexity) with a huge number of reachable vertices.
6. Conclusion and Future Work
In this work, we proposed the single-source RPQ evaluation algorithm, which is based on simultaneous traversal of the input graph and the finite automaton specifying path constraints. The traversal is expressed in terms of operations over matrices and vectors, that allows us to provide highly-parallel implementation based on SuiteSparse:GraphBLAS. Our evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs. While for some queries the proposed algorithm is slower than competitors, we can conclude that our solution is faster for hard queries: it fits with 1 minute time while other solutions do not.
First of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms. For example, rare labels (Koschmieder and Leser, 2012) utilization, or push-pull optimization (Yang et al., 2022) respectively.
While it was shown that multiple source BFS can be expressed in terms of linear algebra (Elekes et al., 2020), there is a room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
Thanks to linear algebra, having a single schema of algorithm one can solve different problems varying semiring-like structures. For example, one can look at variations of BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on used semiring. In the case of RPQ, there are a number of possible outputs and semantics (Angles et al., 2017): reachability, single path, all paths, simple paths, etc. It is an open question, which of them can be expressed without algorithm changes, but by providing other semiring, and which can be expressed with algorithm changes.
Utilization of GPGPUs to evaluate linear algebra based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023). It is necessary to investigate, whether utilization of GPGPU in our algorithm improve performance or not.
Distributed solutions are a way for huge graphs processing (Guo et al., 2021b). Implementation and evaluation of the proposed algorithm in distributed settings, for example, using CombBLAS (Buluç and Gilbert, 2011) that provides distributed linear algebraic routines for graph analysis, is also a task for the future.
Acknowledgements.
This research was supported by the St Petersburg University, grant id 116636233.
Appendix A Proof of Correctness
Theorem A.1 (LA 2-RPQ algorithm correctness).
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
$(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
Label Count
P31 95 749 143
P17 13 994 387
P131 10 609 188
P106 8 533 132
P279 3 128 156
P19 2 800 454
P40 1 419 672
P39 1 302 690
P495 1 272 895
P20 1 127 738
P138 322 330
P551 237 546
Table 2. Number of edges with specific labels used in queries
Query Speedup #Answer
1 Q13442814 ^P31 ?x1 19.09 37 280 216
2 ?x1 (P17 / P31) Q6256 14.49 12 360 733
3 ?x1 (P31* / P279*)+ Q268592 13.34 60 480 400
4 ?x1 ((P106 | P39 | P31) / P279*) Q488383 12.79 76 388 701
5 ?x1 (P31 / P279*) Q191067 11.72 38 410 384
6 ?x1 (P31 / P279*) Q488383 11.10 75 238 748
7 ?x1 (P31 / P279*) Q35120 10.50 90 321 703
8 ?x1 (P31 / P279*) Q13442814 10.44 37 325 506
9 ?x1 (P31 / P279*) Q386724 9.71 55 722 209
10 ?x1 (P31 / P279*) Q7184903 9.53 58 080 849
11 Q10281806 P131* ?x1 0.16 6
12 Q1628288 P131* ?x1 0.17 6
13 Q111 (P31 / P279*) ?x1 0.19 18
14 Q939035 P131* ?x1 0.19 6
15 Q1549583 P131* ?x1 0.20 6
16 Q481075 P131* ?x1 0.20 5
17 Q29873614 P131* ?x1 0.21 4
18 ?x1 P40+ Q748955 0.21 17 699
19 Q28733811 P131* ?x1 0.21 5
20 ?x1 P40+ Q390192 0.23 22 335
Table 3. Top 10 relatively slowest and relatively fastest queries for MDB
Denote state of matrix $M$ after the step $n$ of the loop on lines 7–10 of the algorithm 1 as $M_{n}$ and state of $P$ as $P_{n}$. Introduce auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after the step $n$. $\mathcal{P}$ and $\mathcal{P}_{n}$ are connected with relations $\mathcal{M}_{n}$:
$\mathcal{P}_{n}=\bigcup_{1\leq m\leq n}\mathcal{M}_{m};\;\;\mathcal{P}=\bigcup_{m\in\mathbb{N}}\mathcal{M}_{m}$
The approach is to build a relation between the states of automaton and the vertices of the graph and update it by traversing both automaton and the graph at the same time. As an algorithm invariant after the step $n$ we claim this properties for $\mathcal{M}_{n}$ represented by the matrix $M_{n}$:
$(q,v)\in\mathcal{M}_{n}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{\mathcal{G}}\text{ of length }n\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{\mathcal{N}}\text{ of length }n\text{ in }\mathcal{N}\text{ from }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}\\ \forall m<n\;(q,v)\notin\mathcal{M}_{m}\end{cases}$
Note, that if we continue to loop forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$. It means, it’s enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
Notice, that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty. And $\mathcal{M}_{n+1}\cap\mathcal{P}_{n}=\operatorname{\varnothing}$ for all $n\in\mathbb{N}$. $|\mathcal{P}_{n}|\leq|Q||V|$ and $|\mathcal{M}_{n}|\leq|Q||V|$ for all $n\in\mathbb{N}$. Thus, if $\mathcal{M}_{m}\neq\operatorname{\varnothing}$ for all $m<|Q||V|$ then $\mathcal{M}_{|Q||V|}=\operatorname{\varnothing}$, since $|\mathcal{P}_{|Q||V|-1}|\geq|Q||V|$. This means the algorithm always finishes in maximum of $|Q||V|$ steps.
Obviously, the invariant is held for $n=0$ after the initialization of the matrix $M$ with $M_{0}$. If you consider the paths of length 00 the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
$(q,v)\in\mathcal{M}_{0}\Leftrightarrow q\in Q_{S},v=v_{s}$
Consider evaluating the $n+1$ step of the algorithm. Fix the label $a\in L$. After the step $n$ $M$ represents a relation $\mathcal{M}_{n}$. At first, evaluate the first matrix product $M^{\prime a}_{n+1}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M$. This product represents a relation $\mathcal{M}^{\prime a}_{n+1}\subseteq Q\times V$:
$\mathcal{M}^{\prime a}_{n+1}=\{(q^{\prime},v)\mid(q,q^{\prime})\in\Delta^{a},(q,v)\in\mathcal{M}_{n}\}.$
Evaluate the second matrix product $M^{a}_{n+1}=M^{\prime a}_{n+1}\operatorname{\operatorname{\otimes}}G^{a}=$
$=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M_{n}\operatorname{\operatorname{\otimes}}G^{a}$. Assume $M^{a}_{n+1}$ represents a relation $\mathcal{M}^{a}_{n+1}$, then:
$\mathcal{M}_{n+1}^{a}$ =$=\{(q^{\prime},v^{\prime})\mid(v,v^{\prime})\in E^{a},(q^{\prime},v)\in\mathcal{M}^{\prime a}_{n+1}\}$
=$=\{(q^{\prime},v^{\prime})\mid(v,v^{\prime})\in E^{a},(q,q^{\prime})\in\Delta^{a},(q,v)\in\mathcal{M}_{n}\}.$
Hence, the new relation $\mathcal{M}_{n+1}$ represented by the matrix
$\bigvee_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}M^{a}_{n+1}\langle\neg P\rangle$ can be written as follows:
$\mathcal{M}_{n+1}$ $=\bigcup_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}\{(q,v)\mid(q,v)\in\mathcal{M}^{\prime a}_{n+1},(q,v)\notin\mathcal{P}\}=$
$=\bigcup_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}\left\{(q^{\prime},v^{\prime})\middle|\begin{gathered}(v,v^{\prime})\in E^{a},(q,q^{\prime})\in\Delta^{a}\\ (q,v)\in\mathcal{M}_{n},(q,v)\notin\mathcal{P}\end{gathered}\right\}.$
The updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with $\mathcal{M}^{\prime}$ value is follows form the following properties.
$(q^{\prime},v^{\prime})\in\mathcal{M}_{n+1}\Leftrightarrow\begin{cases}(q,q^{\prime})\in\Delta^{a}\\ (v,v^{\prime})\in E^{a}\\ \end{cases}\text{ for some }a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}.$
Ensure the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$. Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ the paths to $(q,v)$ satisfying the conditions A. Then:
  • $\exists$ 2-path $\pi_{\mathcal{G}}^{\prime}=(e_{1},...,e_{n},(v,v^{\prime}))$ in $\mathcal{G}$ from $v_{s}$ to $v^{\prime}$.
  • $\exists$ path $\pi_{\mathcal{N}}^{\prime}=(\delta_{1},...,\delta_{n},(q,q^{\prime}))$ in $\mathcal{N}$ from $q_{F}\in Q_{F}$ to $q^{\prime}$.
  • $w\cdot a\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}}^{\prime})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}}^{\prime})$ where $w\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})$.
  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, thus, if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
Since, the conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$ the invariant is preserved.
Appendix B Queries Statistics
For detailed analysis of performance we collected top 10 relatively fastest and relatively slowest queries for each system. Results presented in tables 345. Queries represented in SPARQL-like syntax. Speedup means speeding up of the proposed solution relative to respective system. #Answer is a number of reachable vertices. Also we collected total number of edges in the graph with specific labels appeared in collected queries. Results presented in tables 2.
Query Speedup #Answer
1 Q7187 ((P279 / (P279* | P31*)) | (P31/(P279* | P31*))) ?x1 506 771.72 18
2 Q5 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 345 401.17 30
3 Q11629 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 311 512.45 63
4 Q1656682 ((P279 | P31) / (P279* | P31*)) ?x1 306 341.47 7
5 Q5 (P279 / (P279* | P31*)) ?x1 241 437.89 20
6 Q5 ((P279 | P31) / (P279* | P31*)) ?x1 185 795.87 30
7 Q81328 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 127 813.38 18
8 Q4235 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 123 845.99 18
9 Q2 (P31/P279? / P279? / P279? / P279? / P279? / P279?) ?x1 122 229.53 20
10 Q111 (P31 / P279? / P279? / P279? / P279? / P279? / P279?) ?x1 92 952.34 11
11 ?x1 (P17 | P495) Q30 0.03 1 697 681
12 ?x1 (P138 / P106) Q1028181 0.04 8 135
13 ?x1 (P31 / P279) Q18918145 0.05 188 351
14 ?x1 (P19 | P20 | P551) Q90 0.06 50 972
15 ?x1 (P31 / P279) Q16521 0.07 245 925
16 ?x1 (P19 | P20 | P551) Q1741 0.07 43 405
17 ?x1 (P31 / P279*) Q18918145 0.09 217 394
18 ?x1 P31+ Q4830453 0.09 190 623
19 ?x1 (P31 / P279) Q1030034 0.09 82 052
20 ?x1 (P31 / P279) Q486972 0.09 581 652
Table 4. Top 10 relatively slowest and relatively fastest queries for RPQ-Matrix
Query Speedup #Answer
1 ?x1 (P31 / P279*) Q959309 37327.83 1605
2 ?x1 (P31 / P279*) Q385337 30592.48 1980
3 ?x1 (P31/P279*) Q543654 10783.13 7318
4 Q64 (P31/P279*) Q811430 8052.45 115
5 ?x1 P279* Q16707842 2000.93 59
6 ?x1 P279* Q15630906 1664.48 13
7 ?x1 P131* Q168159 1335.28 267
8 ?x1 P279* Q144 606.69 269
9 ?x1 P131* Q108053 598.19 2137
10 ?x1 (P31 / P279*) Q1867183 596.26 1217
11 ?x1 (P138 / P106) Q1028181 0.66 8 135
12 Q257023 P131* ?x1 1.00 6
13 Q326313 P131* ?x1 1.19 6
14 Q204756 P131* ?x1 1.23 5
15 Q29018634 P131* ?x1 1.26 7
16 Q24192592 P131* ?x1 1.26 7
17 Q5964731 P131* ?x1 1.29 6
18 Q328046 P131* ?x1 1.30 6
19 Q2321874 P131* ?x1 1.33 5
20 Q894219 P131* ?x1 1.34 5
Table 5. Top 10 relatively slowest and relatively fastest queries for FalkorDB

### v2

@@
 Abstract.
-A given edge-labelled graph two-way regular path queries (2-RPQs) allow one to use regular languages over labelled edges and inverted edges to constraint paths of interest.
-2-RPQs are (partially) adopted in different real-world graph analysis systems and are a part of the GQL ISO standard.
-But the performance of 2-RPQs on real-world graphs is still a bottleneck for wider adoption.
-A new single-source 2-RPQ algorithm based on linear algebra is proposed.
-Utilization of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries.
-Our implementation demonstrates better performance on average on Wikidata and the respective query log in comparison with MillenniumDB, FalkorDB, and the algorithm of Diego Arroyuelo et al.
-Graph Querying, Regular Path Querying, RPQ, Graph Analysis Algorithms, Sparse Linear Algebra, GraphBLAS
-copyright: acmlicensedjournalyear: 2024conference: ; ; booktitle: ccs: Information systems Graph-based database modelsccs: Information systems Query languages for non-relational enginesccs: Theory of computation Regular languagesccs: Mathematics of computing Paths and connectivity problems
+Two-way regular path queries (2-RPQs) allow one to use regular languages over edges and inverted edges in edge-labelled graph to constrain paths of interest.
+2-RPQs are (partially) adopted in different real-world graph analysis systems and have become a part of the GQL ISO standard.
+However the performance of 2-RPQs on real-world graphs remains a bottleneck for wider adoption.
+Utilisation of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries.
+We propose a new breadth-first-search-based algorithm that leverages linear algebra for evaluating single-source regular path queries.
+We integrate it into the LAGraph graph processing algorithm infrastructure and provide in-depth performance comparison on the large real-world knowledge bases.
+Additionally, we present extensive analysis of its performance across different query types using synthetic data, comparing it with various databases and other linear algebra-based approaches.
 1. Introduction
-Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are formulated in terms of a formal language.
-The language restricts the set of accepted paths: the sentence formed by the labels of a path should be in the language.
+Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are expressed in terms of a formal language.
+The language restricts the set of valid paths: the sequence of labels along a path must form a sentence belonging to the language.
 Regular languages are the most popular class of constraints used as navigational queries in graph databases.
-Queries that use regular languages to specify constraints are called regular path queries or RPQs.
-Being proposed in 1989 by Alberto O.
+Queries that employ regular languages to specify constraints are called regular path queries or RPQs.
+First introduced in 1989 by Alberto O.
 Mendelzon and Peter T.
-Wood (Mendelzon and Wood, 1989), RPQs studying intensively for now.
-Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extend the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b).
-Regular path constraints and related extensions adopted (partially, in some cases) in a wide range of graph analysis systems and respective query languages, such as Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016).
-Moreover, RPQ is a part of ISO standard for a graph query language GQL [^1] ISO/IEC 39075:2024, GQL, https://www.iso.org/standard/76120.html, and core of SPARQL 1.1 RDF query language (201, 2013).
-Nevertheless, having a long history of theoretical and applied research, adoption in real-world solutions, RPQs (and related extensions) are still in the focus of research.
-One of the important direction is an implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real-world cases, and performance-targeted solutions still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024).
-Thus, the development of new performant algorithms for RPQs evaluation is an actual challenge that also pointed out by Angela Bonifati et al in (Bonifati et al., 2018).
-Set of different solutions was proposed to improve RPQ evaluation performance: from special indexes creation (Liu et al., 2021; Arroyuelo et al., 2023b) to utilization of different parallel and distributed models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
-Sparse linear algebra is a promising way to high-performance graph analysis.
-This idea is actively developed by GraphBLAS (Kepner et al., 2016) community [^2] GraphBLAS community web page: https://graphblas.org/.
-Huge amount of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^3] Almost full list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers.
-and respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021).
-Even more recently, it was shown that utilization of sparse linear algebra is a way to a high-performance algorithm for more expressive class of queries, Context-Free Path Queries, or CFPQ (Terekhov et al., 2021).
-But, to our knowledge, there are only a few results on linear algebra based algorithms for RPQ.
-On the one hand, modern graph database FalkorBD444 Sources of FalkorBD on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and utilises linear algebra to analyse graphs.
-This database supports a subset of the Cypher graph query language, including parts of regular constraints.
-But there is no detailed analysis of respective algorithm.
-At the same time, linear algebra based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a).
-The proposed solution directly translates regular expressions to expressions over matrices and vectors.
-Being implemented using sparse matrices and parallel computations, this solution still has performance problems that are evident from the evaluation (Arroyuelo et al., 2023a).
-There are BFS-based strategies for RPQ evaluation (e.g.
-(Koschmieder and Leser, 2012)) [^5] For overview of the primary techniques to evaluate regular path queries, such as relational algebra based and automata based approaches, and ways to optimise it, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018)..
-At the same time, BFS and its variations can be expressed in terms of linear algebra (Davis, 2019a).
-It opens a way to develop a linear algebra based algorithm for RPQ evaluation having BFS-like traversal as the core, and we investigate this way.
-To summarise, in this work we make the following contributions.
+Wood (Mendelzon and Wood, 1989), RPQs have been intensively studied.
+Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extends the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b).
+Regular path constraints and their extensions have been (partially, in some cases) adopted in numerous graph analysis systems and query languages, including Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016).
+Moreover, RPQ is part of the ISO standard for the GQL graph query language (ISO/IEC 39075:2024, 2024), and the core of SPARQL 1.1 RDF query language (201, 2013).
+Despite their long history of theoretical and applied research, as well as real-world adoption, RPQs (and related extensions) remain in the focus of research.
+One of the important directions is the implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real world cases, and performance-targeted solutions are still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024).
+Thus, designing new, efficient algorithms for RPQ evaluation remains an active challenge, as highlighted by Angela Bonifati et al in (Bonifati et al., 2018).
+Various approaches have been proposed to enhance RPQ evaluation performance, ranging from specialized indexing techniques (Liu et al., 2021; Arroyuelo et al., 2023b) to parallel and distributed computing models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
+Sparse linear algebra has emerged as a powerful paradigm for high-performance graph analysis, championed by the GraphBLAS community  (Kepner et al., 2016) [^1] GraphBLAS community web page: https://graphblas.org/.
+A vast number of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^2] An almost complete list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers.
+and the respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021).
+Even more, it has been shown that sparse linear algebra enables a high-performance algorithm for more expressive query classes, such as Context-Free Path Queries (CFPQ) (Terekhov et al., 2021).
+However, to our knowledge, there are only a few studies on linear-algebra-based RPQ algorithms.
+On the one hand, the modern graph database FalkorDB333 Sources of FalkorDB on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and uses linear algebra for graph analysis.
+While FalkorDB supports a subset of the Cypher query language, including some regular constraints, there is no detailed analysis of the respective algorithm.
+At the same time, the linear-algebra-based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a).
+Despite using sparse matrices and parallel computations, this solution exhibits performance limitations, that are evident from the evaluation (Arroyuelo et al., 2023a).
+On the other hand, there are BFS-based strategies for the RPQ evaluation (e.g.
+(Koschmieder and Leser, 2012)) [^4] For an overview of primary RPQ evaluation techniques, including relational algebra and automata-based approaches, and their optimization, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018)..
+Notably, BFS and its variants can be expressed using linear algebra (Davis, 2019a), suggesting an opportunity to develop a linear-algebra-based RPQ algorithm with BFS-like traversal at its core.
+We explore this direction in our work.
+To summarize, in this work we make the following contributions.
   1. (1)
-    A new BFS-inspired algorithm (LA 2-RPQ) for single-source (or single-destination as a symmetric case) 2-RPQ is proposed.
+    A novel BFS-based algorithm (LARPQ) for single-source (or symmetrically, single-destination) 2-RPQ is proposed.
 The algorithm is based on linear algebra: it is expressed in terms of operations over sparse boolean matrices.
-This fact allows one to utilise high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation.
+This fact allows one to utilize high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation.
 The correctness of the proposed algorithm is proven.
   2. (2)
-    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), and with the state-of-the-art graph database MillenniumDB.
-As a result, we show that our solution is comparable with competitors on real-world dataset Wikidata and the respective query log proposed in MillenniumDB path query challenge (Farias et al., 2023).
-While for some queries the proposed algorithm is slower than the competitors, it is faster on average: ×$\times 3.9$ for MillenniumDB, ×$\times 19.6$ for FalkorDB, ×$\times 4.2$ for the algorithm of Diego Arroyuelo et al.
-Moreover, it fits within the 1 minute time limit for all queries, while the other competitors require more time for some queries.
+    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra-based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), with graph databases Blazegraph and state-of-the-art MillenniumDB.
+Experimental results on real-world datasets (Wikidata with query logs from the MillenniumDB path query challenge (Farias et al., 2023) and Yago-2S) demonstrate that our solution achieves competitive performance.
+While occasionally slower on individual queries, our algorithm shows consistent average speedups: 6.8×6.8$\times$6.8 × for the algorithm of Diego Arroyuelo et al, 11.3×11.3$\times$11.3 × for MillenniumDB, 18.9×18.9$\times$18.9 × for FalkorDB, and 16.8×16.8$\times$16.8 × for Blazegraph.
 2. Preliminaries
-In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution.
-2.1. Basics of graph theory
-First we define the graph that we use as a data model.
+In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution and the algorithm proposed by Diego Arroyuelo et al.
+First we define the edge-labeled graph that we use as a data model.
 Definition 2.0 (Edge-labelled graph).
@@
   • $\Delta\subseteq Q\times Q$ is transition relation;
-  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ maps transitions to labels;
+  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ assigns a set of labels (including inverses) to each transition;
   • $Q_{S}\subseteq Q$ is a set of starting states;
   • $Q_{F}\subseteq Q$ is a set of final states.
-The set of languages accepted by 2-NFAs coincides with the set of all RLs over $\Sigma^{\leftrightarrow}$.
+The set of languages accepted by 2-NFAs coincides with the set of all regular languages over $\Sigma^{\leftrightarrow}$.
 Let $[\![\mathcal{N}]\!]=\mathcal{R}$ where $\mathcal{R}$ is the RL accepted by the 2-NFA $\mathcal{N}$.
-In practice special kinds of strings called regular expressions provide a useful facility to define regular languages.
-The algorithms supplying the transformations between regular expressions, RLs and 2-NFAs are out of the scope of this paper.
 Notice that 2-NFA can be seen as a graph $\mathcal{G}_{\mathcal{N}}=\langle Q,\Delta,\Sigma^{\leftrightarrow},\lambda_{\mathcal{N}}\rangle$ equipped with the set of starting states $Q_{S}\subseteq Q$ and the set of final states $Q_{F}\subseteq Q$.
-Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
-Vice versa a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can bee seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$.
+Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
+Conversely, a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can be seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$.
 Such 2-NFA accepts the language:
 $[\![\mathcal{N}_{G}]\!]=\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})$
-Thus, we can treat evaluation of 2-RPQs with fixed set of starting vertices as an intersection of 2-NFAs.
+Thus, we can treat the evaluation of 2-RPQs with a fixed set of starting vertices as an intersection of 2-NFAs.
 Definition 2.0 (2-NFA intersection).
@@
 $[\![\mathcal{N}_{1}\times\mathcal{N}_{2}]\!]=[\![\mathcal{N}_{1}]\!]\cap[\![\mathcal{N}_{2}]\!].$
-We have all the required preliminaries to formally state the problem solved by the 2-RPQ algorithm.
+We have all the necessary preliminaries to formally state the problem solved by the 2-RPQ algorithm.
 Definition 2.0 (Two-way regular path query).
-Recall two-way regular path query (2-RPQ) a triple $\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$ and $v_{s}\in V$ is a starting node.
+Recall two-way regular path query (2-RPQ) a 4-tuple $\langle\mathcal{G},\mathcal{R},V_{s},V_{f}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$, $V_{s}\subseteq V$ is a set of starting nodes, and $V_{f}\subseteq V$ is a set of final nodes.
 Frequent practical cases are queries with fixed starting or final vertex.
-So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ we are interested in an efficient way of evaluating these maps:
-  • Single-source reachability:
+So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},V_{s},V_{f}\rangle$ we are interested in an efficient way of evaluating these maps:
+  • Single-source reachability ($V_{s}=\{v_{s}\},V_{f}=V$ ):
     $[\![\mathcal{Q}]\!]_{SSR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }u\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
-  • Single-destination reachability:
-    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{s}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
+  • Single-destination reachability ($V_{s}=V,V_{f}=\{v_{f}\}$):
+    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{f}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
+For such partial cases we introduce the following short versions: $Q_{SSR}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ and $Q_{SDR}=\langle\mathcal{G},\mathcal{R},v_{f}\rangle$ respectively.
 3. Linear Algebra, Graphs and Relations
-In order to operate with graphs in terms of linear algebra we need to see how algebraic object are connected with set relations and how the relations are connected to graph theory.
+In order to operate with graphs in terms of linear algebra we need to see how algebraic objects are connected with set relations and how the relations are connected to graph theory.
 For two enumerated finite sets $A=\{1,2,...n\}$, $B=\{1,2,...,m\}$ for some $n,m\in\mathbf{N}$ a binary matrix $T$ of size $|A|\times|B|$ can be used to represent a relation $\mathcal{T}\subseteq A\times B$: $T_{ij}=1\Leftrightarrow(i,j)\in\mathcal{T}$.
@@
 Introduce the following operations and constants.
-  • Sum: $A_{n\times m}\operatorname{\operatorname{\oplus}}B_{n\times m}=(a_{ij}\vee b_{ij})_{n\times m}$.
-  • Product: $A_{n\times m}\operatorname{\operatorname{\otimes}}B_{m\times k}=\displaystyle\left(\;\bigvee_{k=1}^{n}a_{ik}\wedge b_{kj}\;\right)_{n\times k}$.
-  • Transposition: $A_{n\times m}^{T}=(a_{ji})_{m\times n}$.
-  • Zero-matrix: $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}=(0)_{n\times k}$.
-  • Mask: $A_{n\times m}\langle C_{n\times m}\rangle=(a_{ij}\wedge c_{ij})_{n\times m}$
-  • Complement: $\neg A=(\neg a_{ij})$
+Definition 3.0.
+Zero matrix is a rectangular matrix $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}$, such that $\operatorname{\operatorname{\mathbf{0}}}_{ij}=0$ for all valid $i$ and $j$.
+Definition 3.0.
+For the Boolean matrix $A_{n\times m}$, the matrix $B_{n\times m}$, such that $B_{ij}=\neg A_{ij}$ is a complement matrix for the matrix $A$. We denote complementation of $A$ as $\neg A$.
+Definition 3.0.
+For the Boolean matrix $A_{n\times m}$, the matrix $B_{m\times n}$, such that $B_{ij}=A_{ji}$ is a transposed matrix for the matrix $A$. We denote transposition of $A$ as $A^{T}$.
+Definition 3.0.
+For the given Boolean matrices $A_{n\times m}$ and $B_{n\times m}$, the sum $A\oplus B$ is a matrix $C_{n\times m}$ such that matrix $C_{ij}=A_{ij}\vee B_{ij}$.
+Definition 3.0.
+For the given Boolean matrices $A_{n\times m}$ and $B_{m\times k}$, the product $A\otimes B$ is a matrix $C_{n\times k}$ such that matrix
+$C_{ij}=\bigvee_{l=1}^{m}A_{il}\wedge B_{lj}.$
+Definition 3.0.
+For the given Boolean matrices $A_{n\times m}$ and $B_{n\times m}$, the masking $A\langle B\rangle$ is a matrix $C_{n\times m}$ such that matrix $C_{ij}=A_{ij}\wedge B_{ij}.$
 Let $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$ be binary relations over sets $S_{1}$, $S_{2}$, $S_{3}$.
@@
            where $\mathcal{A},\mathcal{C}\subseteq S_{1}\times S_{2}$.
-Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an arbitrary graph and $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ an arbitrary 2-NFA.
-Introduce the sets:
+For a given arbitrary graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ and 2-NFA $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ introduce the sets:
   • $E^{a}=\{e\mid e\in E,a\in\lambda^{\leftrightarrow}_{\mathcal{G}}(e)\}$ for all $a\in L^{\leftrightarrow}$.
@@
 This fact strictly leads to the idea of exploiting it for an efficient representation and using sparse matrix operation algorithm implementations.
-Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{-}$ and $G^{a^{-}}=(G^{a})^{T}$.
-4. Single-Source RPQ in Terms of Linear Algebra
+Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{T}$ and $G^{a^{-}}=(G^{a})^{T}$.
+4. BFS-Based Single-Source RPQ in Terms of Linear Algebra
 In this section we describe a single-source linear-algebra based 2-RPQ algorithm and prove its correctness.
 input : $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}}Q_{S},Q_{F}\rangle$, $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $v_{s}\in V$
-output : Matrix $P_{|Q|\times|V|}$
+output : Vector $P^{F}$ of size $1\times|V|$
 1
@@
 3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
-4 $M_{|Q|\times|V|},P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
-5 for $q\in Q_{S}$ do
-       $M_{q{v_{s}}}\leftarrow 1$;
-        // Set the starting relation
-6
-7 end for
-8while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
+4 $P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
+5 $M_{|Q|\times|V|}\leftarrow M$ s.t. $M_{qv}=1$ if $q\in Q_{S}$, $v=v_{s}$, otherwise 00;
+6 $F_{1\times|Q|}\leftarrow F$ s.t. $F_{1}q=1$ if $q\in Q_{F}$, otherwise 00;
+7 while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
        $M\leftarrow\bigoplus_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}((N^{a})^{T}\operatorname{\operatorname{\otimes}}M\operatorname{\operatorname{\otimes}}G^{a})\langle\neg P\rangle$;
         // Update $\mathcal{M}$
-       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$ ;
-        // Accumulate resulting matrix
-9
+8       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$
 9 end while
-11
-return $P$
+10
+return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
 Algorithm 1 Single Source 2-RPQ using linear algebra
-Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA represents query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$.
+Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA that represents the query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$.
 Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be the input graph and $v_{s}\in V$ is a starting node.
@@
 $[\![\mathcal{N}\times\mathcal{N}_{G}]\!]=[\![\mathcal{N}]\!]\cap[\![\mathcal{N}_{\mathcal{G}}]\!]=\mathcal{R}\cap\left(\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})\right).$
-Core of the algorithm is a line 8 that performs one step of traversing two automata simultaneously. Result of the algorithm is a matrix $P_{|Q|\times|V|}$ such that
+Result of the algorithm is a vector $P^{F}_{1\times|V|}$ for which $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$.
+The core of the algorithm is line 8 that performs one step of traversing two automata simultaneously and builds matrix $P_{|Q|\times|V|}$ such that:
 $\begin{cases}P_{q_{F}v}=1\\ q_{F}\in Q_{F}\end{cases}\Leftrightarrow\begin{cases}\exists\pi_{\mathcal{G}}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\mathcal{R}\neq\operatorname{\varnothing}.\end{cases}$
 An example of such a step is presented in figure 1.
-For graph and automaton with labels $\{a,b\}$, we visualise adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation.
-Solid-colored edges should be traversed simultaneously at the step.
-Dashed edges used to represent relation $M$: the edge connects two vertex that are simultaneously reachable.
-Initially, such an edge connects the initial state of the automaton and the starting vertex in the graph.
-After the step (one iteration of the main cycle), there are two such edges.
+For graph and automaton with labels $\{a,b\}$, we visualize adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation.
+The visualization uses solid-colored edges to represent transitions traversed simultaneously at the step, while dashed edges show the reachability relation $M$ — connecting vertices that are simultaneously reachable.
+Initially, $M$ contains just one edge linking the automaton’s initial state to the starting vertex in the graph.
+After one iteration of the main cycle, $M$ contains two edges.
 One of them connects the final state of the automaton with vertex 5 of the graph.
-That means vertex 5 is reachable from the start vertex 3 by the path that farms a word acceptable by the automaton.
-Hence, the set of reachable nodes can be easily obtained from the matrix $P$.
-Let $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ be an evaluated 2-RPQ.
-Then by the main property of the algorithm 1 resulting matrix $P$ is such that
-$v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P_{q_{F}v}=1\text{ for some }q_{F}\in Q_{F}.$
-Procedure 2 computes the explicit representation of the query result.
-Namely, $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$.
-Formally, the following theorem holds.
+This indicates that vertex 5 is reachable from the starting vertex 3 by the path that forms a word acceptable by the automaton.
+The core idea of the algorithm can be summarized in a theorem that is proved using straightforward induction by the length of the paths. Details are provided in appendix A.
 Theorem 4.1 (LA 2-RPQ algorithm correctness).
-The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
+The proposed algorithm, represented in 1, computes the matrix $P$ such that the respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
 $(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
-input : Matrix $P_{|Q|\times|V|}$, set of final states $Q_{F}$
-output : Matrix $P^{F}_{1\times|V|}$
-1
-[…]
-3 for $q\in Q_{F}$ do
-       $F_{1q_{F}}\leftarrow 1$;
-        // Setting up a matrix reducing $P$
-4
-5 end for
-return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
-Algorithm 2 Exporting reachable vertices subprogram
-Proof of this theorem is straightforward induction by the paths length. Details are provided in appendix A.
-Particularly, from the theorem, we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$.
+In particular, from the theorem we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$.
 As far as $q_{F}$ is a final state of the NFA, $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\subseteq[\![\mathcal{N}]\!]$.
@@
 The proposed single-source 2-RPQ algorithm can be used to solve the single-destination problem.
-Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in single destination query.
+Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in the single destination query.
 Denote:
@@
   • $\mathcal{Q}^{-}=\langle\mathcal{G},\mathcal{R}^{-},v_{s}\rangle$.
-The automaton $\mathcal{N}^{-}$ allows traversing paths in the opposite direction.
+The automaton $\mathcal{N}^{-}$ allows one to traverse paths in the opposite direction.
 It holds $\mathcal{R}^{-}=[\![\mathcal{N}^{-}]\!]$.
@@
 Boolean matrix decomposition of $\mathcal{N}^{-}$ can be easily obtained by transposing the matrices in Boolean matrix decomposition of $\mathcal{N}$.
-For performance reasons, it’s necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer non-zero entries.
-Associativity of the Boolean matrix multiplication can be used to perform less calculation.
-This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1:
-$((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
+For performance reasons, it is necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer nonzero entries.
+The associativity of the Boolean matrix multiplication can be used to perform less calculations.
+This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1: $((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
 In general, the first option is preferable due to the fact that 2-NFA for the query can be converted into the corresponding minimal deterministic finite automaton having no more than $|Q|$ non-zero cells in each $\mathcal{N}^{a}$ adjacency matrix due to the automaton determinism.
-This strictly leads to less expensive computation of the product with large $|V|\times|V|$ matrix.
+This strictly leads to a less expensive computation of the product with the large $|V|\times|V|$ matrix.
+When it comes to real-world implementation, it is important to note that it is also possible to implement BFS-based 2-RPQs without using a traversal matrix $M$ as described in algorithm 2.
+The complete description of such algorithm is available in appendix .
+For some sparse linear algebra libraries, it can be more efficient to deal with fewer distinct matrices than to have fewer nonzero entries, especially when there are very few of them.
+Such properties are common for executing simple regular path queries.
+$(q,v)\in\mathcal{P}_{n}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{\mathcal{G}}\text{ of length }\leq n\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{\mathcal{N}}\text{ of length }\leq n\text{ in }\mathcal{N}\text{ from }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
+input : $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}}Q_{S},Q_{F}\rangle$, $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $v_{s}\in V$
+output : Vector $P^{F}$ of size $1\times|V|$
+1
+2$\{N^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{N}$ adjacency matrix;
+3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
+4 $P_{|Q|\times|V|}\leftarrow P$ s.t. $P_{qv}=1$ if $q\in Q_{S}$, $v=v_{s}$, otherwise 00;
+5 $F_{1\times|Q|}\leftarrow F$ s.t. $F_{1}q=1$ if $q\in Q_{F}$, otherwise 00;
+6
+7while $P$ changes do
+       $P\leftarrow P\oplus\bigoplus_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}((N^{a})^{T}\operatorname{\operatorname{\otimes}}P\operatorname{\operatorname{\otimes}}G^{a})$;
+        // Update $\mathcal{P}$
+8
+9 end while
+10
+return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
+Algorithm 2 Single Source 2-RPQ using linear algebra without an extra traversal matrix
+Although it may seem that this algorithm should perform some extra calculations because $P$ contains more entries than $M$ on each step within algorithm 1, this way of implementing the proposed algorithm can be handy.
+Some sparse linear algebra libraries perform better when handling fewer distinct matrices, even if that increases the total number of non-zero entries.
+Such properties are common for executing simple regular path queries.
+4.1. Comparison With Diego Arroyuelo Algorithm
+Diego Arroyuelo et al.
+in (Arroyuelo et al., 2023a) propose another linear algebra-based 2-RPQ evaluation algorithm called RPQ-matrix.
+It directly translates 2-RE to Boolean matrix operations rather than evaluating step-by-step traversal over the graph and the automaton with capability to employ single-source and single-destination 2-RPQs.
+The algorithm consists of the following steps.
+  1. (1)
+    Build an abstract syntax tree (evaluation plan) representing the desired regular expression in which leaves represent labels and other nodes represent operations such as concatenation, Kleene star, and conjunction.
+  2. (2)
+    Match nodes with matrices: each leaf label is matched with an adjacency matrix representing it, and every inner node is matched with a matrix that can be computed based on the child matrices depending on the operation the node represents.
+  3. (3)
+    Optimize the evaluation plan by using the provided source or destination vertex.
+  4. (4)
+    Compute the matrix representing the root element with operations reordering applied to optimize computations.
+We conduct experiments on several graph database management systems and compare them to RPQ-matrix and to the proposed algorithm in order to investigate the efficiency of different linear algebra-based approaches.
 5. Evaluation
-The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) inside the LAGraph (Szárnyas et al., 2021) infrastructure [^6] The source code of the algorithm is available in our GitHub LAGraph fork: https://github.com/SparseLinearAlgebra/LAGraph/tree/rpq.
-Both sparse matrices and their transpositions are loaded in-memory due to the fact we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
-Whereas we’re focusing on an efficient way of evaluating RPQs in-memory, there aren’t many candidates to compare with.
+The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) within the LAGraph (Szárnyas et al., 2021) infrastructure [^5] Regular Path Query algorithm in LAGraph repository: https://github.com/GraphBLAS/LAGraph/blob/stable/experimental/algorithm/LAGraph_RegularPathQuery.c.
+Both sparse matrices and their transpositions are loaded in memory due to the fact that we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
+Whereas we are focusing on an efficient way of evaluating RPQs in memory, there are not many candidates to compare.
 Popular database management systems primarily focus on general availability and ensure availability of the concurrent access at the same time the suggested algorithm implementation is suited only for solving the reachability problem.
-We’ve selected the following systems as demonstrating the best performance and being the most related to our use case.
-MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system supporting a synthetic way to carry out calculations without using the disk storage by caching the query data.
-It demonstrates state-of-the-art performance on evaluating regular path queries.
-FalkorDB (ex.
-RedisGraph (Cailliau et al., 2019)) is an in-memory property-graph database also employing SuiteSparse:GraphBLAS for query evaluation.
-However, it uses OpenCypher modification and supports only a subset of regular path queries.
-To deal with this, we’ve carried out the measurements of cypher-compatible and non-cypher-compatible queries separately (e.g., it is impossible to evaluate repeated-path queries in the form of $(ab)^{*}$ in FalkorDB).
-RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the in-memory linear-algebra based RPQ evaluation algorithm.
+We have selected the following systems as the most effective and the most related to our use case.
+RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the linear algebra-based RPQ evaluation algorithm.
 The original work introduces a few variations of the adjacency matrix representation: $k^{2}$-trees offering less memory consumption and CSR/CSC formats providing better performance.
-We’ve chosen the last one to compare with since we’re aiming to compare the performance.
+We have chosen the last one to compare since we are aiming to compare the performance.
+RPQ-matrix (GrB) [^6] SuiteSparse:GraphBLAS-based implementation of RPQ-matrix: https://github.com/suvorovrain/rpq-matrix/tree/gbmod is an adapted version of Diego Arroyuelo et al.
+RPQ-matrix algorithm where matrix representations and operations are substituted with SuiteSparse:GraphBLAS equivalents in order to analyse performance impact of basic primitives implementation.
+MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system with RDF-model and SPARQL support.
+It supports a synthetic way to carry out calculations without using disk storage by caching query data.
+MillenniumDB demonstrates state-of-the-art performance on evaluating regular path queries.
+FalkorDB (previously RedisGraph (Cailliau et al., 2019)) is an in-memory property graph database that also employs SuiteSparse:GraphBLAS for query evaluation.
+However, it uses OpenCypher modification and supports only a subset of regular path queries (e.g., it is impossible to evaluate repeated-path queries in the form of $(a\;b)^{*}$).
+To deal with this, we have carried out the measurements of cypher-compatible and non-cypher-compatible queries separately.
+Blazegraph (Vrgoc et al., 2021) is another graph-oriented database management system using RDF data model and SPARQL query language.
+It is used by the Wikidata project and is based on more classical B-trees.
 All experiments are conducted on a work station with Ryzen 9 7900X 4.7 GHz 12-core, 128 Gb of DDR5 RAM and running Ubuntu 22.04.
+5.1. Implementation Details
+As mentioned above, it is important for better performance to take into account that there are two different possible implementations of the BFS-based 2-RPQ algorithm: one involving less different matrices and one with less dense matrices.
+For SuiteSparse:GraphBLAS, the first approach turns out to be faster if the number of entries in matrices is small, and the latter is better for denser matrices.
+The most straightforward way to combine these two approaches to achieve better performance is to switch from the first approach to the second one during the traversal if the resulting matrix starts having some constant amount of non-zero entries.
+This constant can be empirically determined for the graph [^7] For the studied datasets the most suitable value is $100$..
+This allows the BFS-based algorithm to provide strong performance on simple queries with very few answers and on analytical queries involving a lot of resulting vertices at the same time.
 5.2. Dataset Description
+To compare performance of different approaches we start from evaluating benchmarks on large real-world datasets.
 For algorithm evaluation, we choose the Wikidata dataset from the snapshot provided in terms of MillenniumDB path query challenge (Farias et al., 2023).
 The resource contains both the graph and the set of 660 different 2-RPQs in SPARQL format taken from the Wikidata query log.
-For systems that do not support RDF format, the dataset was converted to an edge-labelled graph.
-The processed graph has 610 402 396610402396610\,402\,396610 402 396 edges, 91 609 2549160925491\,609\,25491 609 254 vertices and 1 39513951\,3951 395 distinct labels.
-SPARQL path queries has been converted to the corresponding minimized DFAs.
-For evaluating on FalkorDB the supported subset of regular path queries have been converted into Cypher queries of form MATCH ...
-COUNT (DISTINCT ...).
-The original dataset contains queries from all to all vertices and queries involving non-existing entities.
-They’ve been filtered out, resulting in 578 queries remaining.
-5.2. Evaluation Results
+The second dataset is Yago-2S evaluated with 7 different complex queries taken from (Abul-Basher et al., 2017).
+We also want to compare different linear algebra-based approaches and determine the best of them for different query kinds.
+Real-world datasets are not suitable for it due to complex graph topology.
+Instead, we employ the synthetic RPQBench dataset generator (Wang et al., 2025) for controlled performance evaluation across query categories.
+The structure of the graph ensures reproducible and predictable results when similar queries are executed.
+The original generator produces arbitrary sized RDF datasets and offers $10$ different query kinds without starting or final nodes specified in SPARQL format.
+We generate a graph and supply these query kinds with randomly generated source and destination vertices.
+For systems that do not support the RDF format, the datasets have been converted to an edge-labeled graph.
+Original SPARQL path queries have been deprefixed, converted to the corresponding minimized DFAs.
+Queries have been converted to Cypher queries of the form MATCH ...
+COUNT (DISTINCT ...) when it is possible.
+Queries without starting or final vertex, broken queries involving missing entities have been removed.
+Final dataset statistics can be summarised as follows.
+Wikidata
+  • $610$ million edges, $91$ million vertices, $1400$ distinct labels.
+  • $578$ queries filtered out from $660$ from the query dump.
+Yago-2S
+  • $610$ million edges, $91$ million vertices, $1400$ distinct labels.
+  • $7$ complex queries taken from (Abul-Basher et al., 2017).
+RPQBench
+  • Synthetic dataset, $610$ million edges, $91$ million vertices, $1400$ distinct labels.
+  • $20$ different query kinds supplied $1000$ randomly generated source/destination vertices.
+  • Trivial queries are filtered out (e.g. $a^{*}$ has $\geq 1$ answer).
+5.3. Evaluation Scenario
+((a)) Queries which are supported and succeeded on all of the competitors
+((b)) Queries which are not supported in Cypher or timed out on FalkorDB or Blazegraph
 Figure 2. Wikidata dataset per-query evaluation time
-LA 2-RPQ MilenniumDB RPQ-Matrix FalkorDB
-Mean C (ms) 83 327 348 1 625
-Median C (ms) 2.5 4.4 4.2 18.8
-Mean speedup C 1 3.9 4.2 19.6
-Median speedup C 1 1.8 1.7 7.5
-Mean NC (ms) 374 1 345 7 108
-Median NC (ms) 37.4 69.8 269.6
-Mean speedup NC 1 3.6 19.0
-Median speedup NC 1 1.9 7.2
-Memory (Gb) 18.1 6.9 50.0
-BPT 31.9 12.1 85.8
-Timeouts 0 1 15 53
 Table 1.
-Query execution results.
-Results are provided for cypher-compatible (C) and non-cypher-compatible (NC) queries separately.
+Wikidata query execution results.
+The results are supplied for simple and for complex (C) queries separately.
+For in-memory algorithms the memory consumption for whole dataset and bytes-per-triple values (BPT) are presented.
 Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
-We measure evaluation time per query with a timeout of 1 minute. All of the data required for query evaluation is pre-loaded into the memory. Queries are executed one by one.
-We run MillenniumDB [^7] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 in a SPARQL mode with the maximum amount of CPU cores available.
-The query dataset is executed twice.
-The first run isn’t measured since it makes the database load the data into the memory.
-The query execution time is calculated as the internal optimizer plus execution time excluding the parser.
-We execute FalkorDB v4.2.0888FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 querying through the python wrapper.
-The database is started without CPU core limitations and without query cache.
-We build an index over the vertices to be able to select the starting or destination vertex using it.
-The query execution time is based on the internal timer.
-The RPQ-matrix algorithm is executed in CSR/CSC mode without core and memory consumption limitations.
-The consumed time is measured with the internal timer.
-The solution has been built out of the project repository [^9] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606.
-Results of the measurements, both evaluation time and relative speedup, are provided in figure 2.
-Only queries expressible in Cypher are presented.
-The purple lines represent mean values, and the green lines represent medians.
-Note that relative speedup is represented as pointwise speedup for each query.
-So, mean and median calculated for per-query speedup.
-Numeric values of the mean and the median, the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are available within the table 1.
-Queries that have failed on any of the systems with timeouts are excluded from the statistics.
-The number of time outed queries shown in table 1.
-We can see that while in median, speeding up for our solution is between 1.7 and 7.5, in mean we improve performance more than 19 times (relative to the most closest solutions RPQ-Matrix and FalkorDB).
-That means the proposed solution significantly outperforms other linear algebra based solutions on some queries (up to ×$\times 10^{5}$) while demonstrating comparable (and better in average) performance in the most cases.
-On the other hand, RPQ-Matrix demonstrates better performance on big number of queries.
-It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based solutions (proposed and RPQ-Matrix) to take the best from both of them.
-Note that MillenniumDB is pretty stable: there is neither outstanding wins nor losses in performance.
-At the same time, the proposed solution is slightly better in average even though MillenniumDB caches query evaluation data, but our solution does not.
-Memory consumption is presented on table 1.
-We don’t specify MillenniumDB memory consumption due to the fact it can process query in-memory only by caching data and doesn’t provide ways to preload whole dataset from the disk storage.
-RPQ-Matrix demonstrates the lowest memory usage.
-Memory consumption of the proposed solution is comparable with competitors.
-We collected the top 10 queries with the worst and best speedup relative to all competitors.
-Collected data presented in tables 345.
-We can see that the weakness of our solution are relatively simple queries with a relatively small number of reachable vertices, like Q P131* ?x1 with less than 100 reachable vertices, or ?x1 (P138 / P106) Q with less than a million reachable vertices, where Q is a fixed vertex, and ?x1 is a variable.
-The forte of the proposed solution is a complex query (in terms of regexp complexity) with a huge number of reachable vertices.
+LARPQ RPQ-matrix RPQ-matrix (GrB) MillenniumDB FalkorDB Blazegraph
+Total, ms 24 403 166 882 34 891 276 527 460 606 403 738
+Mean, ms 49.8 340.6 71.2 564.3 940.0 824.0
+Median, ms 2.0 3.3 1.1 13.7 17.6 128.0
+Mean speedup 1.00 0.15 0.70 0.09 0.05 0.06
+Median speedup 1.00 0.60 1.74 0.15 0.11 0.02
+Total C, ms 35 061 508 378 288 766 639 193
+Mean C, ms 427.6 6 199.7 3 521.5 7 795.0
+Median C, ms 169.1 372.9 142.9 3 105.6
+Mean speedup C 1.00 0.07 0.12 0.05
+Median speedup C 1.00 0.45 1.18 0.05
+Memory, Gb 9.2 6.9 9.2
+BPT 16.3 12.1 16.3
+Table 2. Yago-2S query execution results. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
+LARPQ RPQ-matrix
+RPQ-matrix
+MillenniumDB Blazegraph
+Total, ms 377 174 202 7 025 24883
+Mean, ms 54 25 29 1 004 3555
+Median, ms 69 25 38 985 3808
+Mean speedup 1.00 2.17 1.87 0.05 0.02
+Median speed up 1.00 2.71 1.81 0.07 0.02
+Memory, Gb 0.5 0.4 0.5
+BPT 11.3 9.8 11.3
+We measure query evaluation time with a 1-minute timeout, preloading all required data into memory, so data preprocessing time is not included. Queries execute sequentially in isolation.
+RPQ-Matrix [^8] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606 configuration.
+  • CSR/CSC mode as the most performant one.
+  • No CPU/cores limitations.
+  • No memory limitations.
+  • Timing: internal timer.
+RPQ-matrix (GrB) [^9] RPQ-matrix implemented using SuiteSparse:GraphBLAS repository: https://github.com/suvorovrain/rpq-matrix/tree/gbmod configuration.
+  • Pre-caching CSR/CSC matrices.
+  • No CPU/cores limitations.
+  • No memory limitations.
+  • Timing: internal timer.
+MillenniumDB [^10] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 configuration.
+  • SPARQL mode.
+  • Maximum CPU cores.
+  • Two execution runs (first warm-up excluded).
+  • Timing: internal timer for query optimization + execution (parser excluded).
+FalkorDB (v4.2.0) [^11] FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 configuration.
+  • Accessed via Python wrapper.
+  • No CPU/cores limitations.
+  • No query caching.
+  • Vertices index for efficient start/end vertex selection.
+  • Timing: internal database timer.
+Blazegraph (2.1.5) [^12] Blazegraph v2.1.5: https://github.com/blazegraph/database/tree/BLAZEGRAPH_RELEASE_2_1_5 configuration.
+  • Accessed via Python wrapper.
+  • No CPU/cores limitations.
+  • No memory limitations.
+  • Timing: external timer.
+  • Queries are executed one by one with extra heatup to avoid out-of-memory.
+All benchmarking scripts, configurations, and instructions are available in GitHub repository [^13] Benchmark evaluation repository: https://github.com/SparseLinearAlgebra/la-rpq.
+5.4. Real-World Graph Querying Results
+Resulting per-query evaluation time is provided in Figure 2(a) for queries that are supported on all of the competitors.
+Remaining complex queries ended up with a timeout on slower competitors or not expressible in Cypher are represented separately in Figure 2(b).
+The dotted lines represent means and the straight lines represent medians.
+These numeric values of and total execution time are available in the table 1 for the relatively simple and complex queries (C) separately.
+For linear algebra-based competitors the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are provided.
+For simple queries, LARPQ demonstrates the best mean with a speedup of 1.7×1.7$\times$1.7 × to 18.9×18.9$\times$18.9 ×.
+However, RPQ-matrix implemented with SuiteSparse:GraphBLAS demonstrates the best median time that is better 1.7×1.7$\times$1.7 × than that of LARPQ.
+It means that RPQ-matrix evaluates some of the queries faster whereas the BFS-based algorithm works better in general cases.
+Our hypothesis is that the proposed algorithm does not utilise the sparsity of some edge kinds enough.
+Both linear algebra-based approaches demonstrate better time in mean an in median in comparison to all of database management systems.
+For complex queries, relations between different competitors are the same except that LARPQ mean is drastically lower than other competitors being 14.5×14.5$\times$14.5 ×, 8.3×8.3$\times$8.3 ×, 18.3×18.3$\times$18.3 × less than those of the competitors.
+It means the proposed algorithm is capable of executing the most complex out of the complex algorithm faster than other approaches.
+The results for the Yago-2S dataset are presented in table 2. FalkorDB is excluded since every of the $7$ queries have resulted with a timeout.
+Both RPQ-matrix algorithms demonstrate better time than the proposed LARPQ algorithm.
+It is likely due to the structure of the query.
+All of them use $a\;b\;c^{+}\;d^{+}$ patterns.
+Iterative structure of the BFS-based approach does not utilize efficient evaluation order that is important for such long queries of simple operations.
+It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based-solutions (proposed and RPQ-Matrix) to take the best from both of them.
+5.5. Synthetic Graph Querying Results
+We perform a detailed comparison of competitors for executing different query kinds by running a synthetic RPQBench.
+Its evaluation results are available in the table 3.
+Each row represents distinct query kind.
+Total execution time of the randomly generated queries are presented in seconds for each competitor separately.
+Due to the lower overall performance demonstrated on the real-world datasets, MillenniumDB, FalkorDB, and Blazegraph are excluded from subsequent comparisons.
+We also provide edge statistics for each label kind in table 4.
+Table 3. RPQBench dataset evaluation time of queries with randomly generated sources and destinations per each query kind in seconds
+Query pattern, single-source
+(S) or single-destination (D)
+LARPQ RPQ-matrix
+RPQ-matrix
+1 $a\;b\;c$, (S) 51 83 11
+2 $a\;b\;c$, (D) 32 11 12
+3 $(a\;b\;c)\mid(c\;d\;d)$, (S) 59 89 24
+4 $(a\;b\;c)\mid(c\;d\;d)$, (D) 47 14 22
+5 $d^{*}$, (S) 22 21 37
+6 $d^{*}$, (D) 19 19 23
+7 $d^{*}\;e$, (S) 21 12 18
+8 $d^{*}\;e$, (D) 6 3 5
+9 $d^{+}$, (S) 23 14 29
+10 $d^{+}$, (D) 18 14 20
+11 $(a\;b)^{*}$, (S) 1 350 30 156 4 904
+12 $(a\;b)^{*}$, (D) 3 619 15 706 2 978
+13 $f\;g\;(d\mid e)$, (S) 92 700 5 917 7 745
+14 $f\;g\;(d\mid e)$, (D) 2 193 301 486
+15 $f\;g\;(d\mid e)^{*}$, (S) 149 644 2 870 709 2 053 090
+16 $f\;g\;(d\mid e)^{*}$, (D) 6 167 810 1 053
+17 $(c\mid g)\;(d\mid e)$, (S) 36 3 7
+18 $(c\mid g)\;(d\mid e)$, (D) 16 173 017 41 713
+19 $(c\mid g)\;(d\mid e)^{*}$, (S) 930 55 232
+20 $(c\mid g)\;(d\mid e)^{*}$, (D) 99 955 891 229 035
+Table 4. RPQBench edge stats per label
+Edge label Count Edge label Count
+$a$ 343 660 $e$ 36
+$b$ 4 209 447 $f$ 4 928 456
+$c$ 114 742 222 $g$ 223 656
+$d$ 186
+As it is observed, the execution time of the proposed BFS-based algorithm for simple queries such as $d^{*}$, $d^{+}$, and $d^{*}e$ is quite similar to that of other linear algebra-based implementations.
+RPQ-matrix ×$\times 2$ speedups are likely to happen due to the CSR/CSC matrix implementation since RPQ-matrix (GrB) demonstrates evaluation time close to LARPQ.
+[…]
+For the remaining queries, such as 1–4, the proposed algorithm demonstrates a relative slowdown of 2×2$\times$2 × to 4×4$\times$4 ×, since it does not utilize an efficient evaluation order.
+For queries 14, 16, 17, and 19, LARPQ is 4×4$\times$4 × to 17×17$\times$17 × slower than the competitors because it does not exploit the fact that edges with labels $d$ or $e$ are very rare.
+Conclusion about various linear algebra-based algorithms might be summarized as follows.
+  • LARPQ is a better choice for queries that involve compound operations over labels corresponding to many edges and not having source/destination vertices close to them.
+  • LARPQ algorithm tends to be more stable whereas RPQ-matrix might demonstrate drastic slowdown for some kinds of complex queries.
+  • Both algorithms are efficient enough for simple queries.
+  • RPQ-matrix is a better choice if the query contains rare labels, long concatenations or disjunctions allowing to perform optimizations.
 6. Conclusion and Future Work
-In this work, we proposed the single-source RPQ evaluation algorithm, which is based on simultaneous traversal of the input graph and the finite automaton specifying path constraints.
-The traversal is expressed in terms of operations over matrices and vectors, that allows us to provide highly-parallel implementation based on SuiteSparse:GraphBLAS.
-Our evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs.
+In this work, we proposed the single-source RPQ evaluation algorithm, which is based on the simultaneous traversal of the input graph and the finite automaton specifying path constraints.
+The traversal is expressed in terms of operations over matrices and vectors, which allows us to provide a highly parallel implementation based on SuiteSparse:GraphBLAS.
+Our experimental evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs.
 While for some queries the proposed algorithm is slower than competitors, we can conclude that our solution is faster for hard queries: it fits with 1 minute time while other solutions do not.
-First of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms.
+Our results demonstrate that both algorithmic design and implementation details of underlying linear algebra primitives significantly impact 2-RPQ performance.
+While average performance metrics provide straightforward comparisons, specific cases require deeper analysis to identify each algorithm’s strengths under different conditions; to understand implementation trade-offs; to guide development of robust universal solutions.
+Furthermore, this analysis provides valuable insights for enhancing the GraphBLAS API by identifying critical functionality required for efficient RPQ evaluation algorithms.
+The performance characteristics we observed highlight specific linear algebra primitives that most significantly impact query processing efficiency, suggesting potential directions for API optimization.
+Regarding our new algorithm, first of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms.
 For example, rare labels (Koschmieder and Leser, 2012) utilization, or push-pull optimization (Yang et al., 2022) respectively.
-While it was shown that multiple source BFS can be expressed in terms of linear algebra (Elekes et al., 2020), there is a room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
+Although multiple source BFS has been shown to be expressed in terms of linear algebra (Elekes et al., 2020), there is room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
 Thanks to linear algebra, having a single schema of algorithm one can solve different problems varying semiring-like structures.
-For example, one can look at variations of BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on used semiring.
+For example, one can look at variations of the BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on the used semiring.
 In the case of RPQ, there are a number of possible outputs and semantics (Angles et al., 2017): reachability, single path, all paths, simple paths, etc.
-It is an open question, which of them can be expressed without algorithm changes, but by providing other semiring, and which can be expressed with algorithm changes.
-Utilization of GPGPUs to evaluate linear algebra based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023).
-It is necessary to investigate, whether utilization of GPGPU in our algorithm improve performance or not.
-Distributed solutions are a way for huge graphs processing (Guo et al., 2021b).
+It is an open question, which of them can be expressed without algorithm changes, but by providing other semirings, and which can be expressed with algorithm changes.
+Utilization of GPGPUs to evaluate linear algebra-based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023).
+It is necessary to investigate, whether utilization of GPGPU in our algorithm improves performance or not.
+Distributed solutions are a way to process graph processing (Guo et al., 2021b).
 Implementation and evaluation of the proposed algorithm in distributed settings, for example, using CombBLAS (Buluç and Gilbert, 2011) that provides distributed linear algebraic routines for graph analysis, is also a task for the future.
 Acknowledgements.
-This research was supported by the St Petersburg University, grant id 116636233.
+This research has been supported by the St. Petersburg State University, grant id 116636233. Paper submission has been sponsored by Tarantool DBMS team, VK Tech LLC.
 Appendix A Proof of Correctness
 Theorem A.1 (LA 2-RPQ algorithm correctness).
-The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
+The proposed algorithm, represented in 1, computes the matrix $P$ such that the respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
 $(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
-Label Count
-P31 95 749 143
-P17 13 994 387
-P131 10 609 188
-P106 8 533 132
-P279 3 128 156
-P19 2 800 454
-P40 1 419 672
-P39 1 302 690
-P495 1 272 895
-P20 1 127 738
-P138 322 330
-P551 237 546
-Table 2. Number of edges with specific labels used in queries
-Query Speedup #Answer
-1 Q13442814 ^P31 ?x1 19.09 37 280 216
-2 ?x1 (P17 / P31) Q6256 14.49 12 360 733
-3 ?x1 (P31* / P279*)+ Q268592 13.34 60 480 400
-4 ?x1 ((P106 | P39 | P31) / P279*) Q488383 12.79 76 388 701
-5 ?x1 (P31 / P279*) Q191067 11.72 38 410 384
-6 ?x1 (P31 / P279*) Q488383 11.10 75 238 748
-7 ?x1 (P31 / P279*) Q35120 10.50 90 321 703
-8 ?x1 (P31 / P279*) Q13442814 10.44 37 325 506
-9 ?x1 (P31 / P279*) Q386724 9.71 55 722 209
-10 ?x1 (P31 / P279*) Q7184903 9.53 58 080 849
-11 Q10281806 P131* ?x1 0.16 6
-12 Q1628288 P131* ?x1 0.17 6
-13 Q111 (P31 / P279*) ?x1 0.19 18
-14 Q939035 P131* ?x1 0.19 6
-15 Q1549583 P131* ?x1 0.20 6
-16 Q481075 P131* ?x1 0.20 5
-17 Q29873614 P131* ?x1 0.21 4
-18 ?x1 P40+ Q748955 0.21 17 699
-19 Q28733811 P131* ?x1 0.21 5
-20 ?x1 P40+ Q390192 0.23 22 335
-Table 3. Top 10 relatively slowest and relatively fastest queries for MDB
-Denote state of matrix $M$ after the step $n$ of the loop on lines 7–10 of the algorithm 1 as $M_{n}$ and state of $P$ as $P_{n}$.
-Introduce auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after the step $n$.
+Denote the state of matrix $M$ after the step $n$ of the loop on lines 7–10 of algorithm 1 as $M_{n}$ and the state of $P$ as $P_{n}$.
+Introduce the auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after step $n$.
 $\mathcal{P}$ and $\mathcal{P}_{n}$ are connected with relations $\mathcal{M}_{n}$:
 $\mathcal{P}_{n}=\bigcup_{1\leq m\leq n}\mathcal{M}_{m};\;\;\mathcal{P}=\bigcup_{m\in\mathbb{N}}\mathcal{M}_{m}$
-The approach is to build a relation between the states of automaton and the vertices of the graph and update it by traversing both automaton and the graph at the same time.
+The approach is to build a relation between the automaton states and the vertices of the graph and update it by traversing both automaton and graph at the same time.
 As an algorithm invariant after the step $n$ we claim this properties for $\mathcal{M}_{n}$ represented by the matrix $M_{n}$:
 $(q,v)\in\mathcal{M}_{n}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{\mathcal{G}}\text{ of length }n\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{\mathcal{N}}\text{ of length }n\text{ in }\mathcal{N}\text{ from }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}\\ \forall m<n\;(q,v)\notin\mathcal{M}_{m}\end{cases}$
-Note, that if we continue to loop forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$.
-It means, it’s enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
-Notice, that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty.
+Note that if we continue looping forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$.
+This means that it is enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
+Notice that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty.
 And $\mathcal{M}_{n+1}\cap\mathcal{P}_{n}=\operatorname{\varnothing}$ for all $n\in\mathbb{N}$.
@@
 Thus, if $\mathcal{M}_{m}\neq\operatorname{\varnothing}$ for all $m<|Q||V|$ then $\mathcal{M}_{|Q||V|}=\operatorname{\varnothing}$, since $|\mathcal{P}_{|Q||V|-1}|\geq|Q||V|$.
-This means the algorithm always finishes in maximum of $|Q||V|$ steps.
-Obviously, the invariant is held for $n=0$ after the initialization of the matrix $M$ with $M_{0}$.
-If you consider the paths of length 00 the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
+This means that the algorithm always finishes in a maximum of $|Q||V|$ steps.
+Obviously, the invariant holds for $n=0$ after initializing the matrix $M$ with $M_{0}$.
+If you consider the paths of length 00, the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
 $(q,v)\in\mathcal{M}_{0}\Leftrightarrow q\in Q_{S},v=v_{s}$
@@
 Fix the label $a\in L$.
-After the step $n$ $M$ represents a relation $\mathcal{M}_{n}$.
+After step $n$, $M$ represents a relation $\mathcal{M}_{n}$.
 At first, evaluate the first matrix product $M^{\prime a}_{n+1}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M$.
@@
 $=\bigcup_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}\left\{(q^{\prime},v^{\prime})\middle|\begin{gathered}(v,v^{\prime})\in E^{a},(q,q^{\prime})\in\Delta^{a}\\ (q,v)\in\mathcal{M}_{n},(q,v)\notin\mathcal{P}\end{gathered}\right\}.$
-The updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with $\mathcal{M}^{\prime}$ value is follows form the following properties.
+Updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with the $\mathcal{M}^{\prime}$ value derives the following properties.
 $(q^{\prime},v^{\prime})\in\mathcal{M}_{n+1}\Leftrightarrow\begin{cases}(q,q^{\prime})\in\Delta^{a}\\ (v,v^{\prime})\in E^{a}\\ \end{cases}\text{ for some }a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}.$
-Ensure the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$.
-Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ the paths to $(q,v)$ satisfying the conditions A.
+Ensure that the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$.
+Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ be the paths to $(q,v)$ satisfying conditions A.
 Then:
@@
   • $w\cdot a\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}}^{\prime})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}}^{\prime})$ where $w\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})$.
-  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, thus, if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
-Since, the conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$ the invariant is preserved.
-Appendix B Queries Statistics
-For detailed analysis of performance we collected top 10 relatively fastest and relatively slowest queries for each system.
-Results presented in tables 345.
-Queries represented in SPARQL-like syntax.
-Speedup means speeding up of the proposed solution relative to respective system.
-#Answer is a number of reachable vertices.
-Also we collected total number of edges in the graph with specific labels appeared in collected queries.
-Results presented in tables 2.
-Query Speedup #Answer
-1 Q7187 ((P279 / (P279* | P31*)) | (P31/(P279* | P31*))) ?x1 506 771.72 18
-2 Q5 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 345 401.17 30
-3 Q11629 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 311 512.45 63
-4 Q1656682 ((P279 | P31) / (P279* | P31*)) ?x1 306 341.47 7
-5 Q5 (P279 / (P279* | P31*)) ?x1 241 437.89 20
-6 Q5 ((P279 | P31) / (P279* | P31*)) ?x1 185 795.87 30
-7 Q81328 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 127 813.38 18
-8 Q4235 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 123 845.99 18
-9 Q2 (P31/P279? / P279? / P279? / P279? / P279? / P279?) ?x1 122 229.53 20
-10 Q111 (P31 / P279? / P279? / P279? / P279? / P279? / P279?) ?x1 92 952.34 11
-11 ?x1 (P17 | P495) Q30 0.03 1 697 681
-12 ?x1 (P138 / P106) Q1028181 0.04 8 135
-13 ?x1 (P31 / P279) Q18918145 0.05 188 351
-14 ?x1 (P19 | P20 | P551) Q90 0.06 50 972
-15 ?x1 (P31 / P279) Q16521 0.07 245 925
-16 ?x1 (P19 | P20 | P551) Q1741 0.07 43 405
-17 ?x1 (P31 / P279*) Q18918145 0.09 217 394
-18 ?x1 P31+ Q4830453 0.09 190 623
-19 ?x1 (P31 / P279) Q1030034 0.09 82 052
-20 ?x1 (P31 / P279) Q486972 0.09 581 652
-Table 4. Top 10 relatively slowest and relatively fastest queries for RPQ-Matrix
-Query Speedup #Answer
-1 ?x1 (P31 / P279*) Q959309 37327.83 1605
-2 ?x1 (P31 / P279*) Q385337 30592.48 1980
-3 ?x1 (P31/P279*) Q543654 10783.13 7318
-4 Q64 (P31/P279*) Q811430 8052.45 115
-5 ?x1 P279* Q16707842 2000.93 59
-6 ?x1 P279* Q15630906 1664.48 13
-7 ?x1 P131* Q168159 1335.28 267
-8 ?x1 P279* Q144 606.69 269
-9 ?x1 P131* Q108053 598.19 2137
-10 ?x1 (P31 / P279*) Q1867183 596.26 1217
-11 ?x1 (P138 / P106) Q1028181 0.66 8 135
-12 Q257023 P131* ?x1 1.00 6
-13 Q326313 P131* ?x1 1.19 6
-14 Q204756 P131* ?x1 1.23 5
-15 Q29018634 P131* ?x1 1.26 7
-16 Q24192592 P131* ?x1 1.26 7
-17 Q5964731 P131* ?x1 1.29 6
-18 Q328046 P131* ?x1 1.30 6
-19 Q2321874 P131* ?x1 1.33 5
-20 Q894219 P131* ?x1 1.34 5
-Table 5. Top 10 relatively slowest and relatively fastest queries for FalkorDB
+  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, so if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
+Since conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$, the invariant is preserved.